# Engineering Anderson Localization in Arbitrary Dimensions with Interacting Quasiperiodic Kicked Bosons

### v1

@@ 4 Conclusion
Beyond its conceptual simplicity, this approach provides a powerful tool to explore Anderson localization and quantum criticality in regimes that are otherwise experimentally challenging to reach. Future directions include the development of experimental implementations using ultracold atomic gases or photonic systems, where both interactions and quasiperiodic driving can be precisely controlled. From a theoretical standpoint, our framework opens the way to investigate other symmetry classes, such as the unitary or symplectic ones, by appropriately modifying the driving protocol or introducing synthetic gauge fields. It also provides a natural platform to study properties beyond averaged quantities, including multifractality and nonergodic dynamics at criticality.
@@ Appendix A Calculation of the matrix elements in the Bethe basis
In this section, we recall the expressions of the eigenstates of the Hamiltonian (2) from the main text for the case of $N=2$ identical bosons [LiebLiniger] and give the explicit expression of the matrix elements of the Floquet operator in this basis. Following the Bethe Ansatz, the wavefunction in each position sector is expressed as a superposition of plane waves with fixed energy. In the fundamental position sector $x_{1}<x_{2}$, the Bethe wavefunction takes the form
@@
where the positions $x_{j}$ are expressed in units of $\ell=L/(2\pi).$ The non-normalized amplitudes $A_{ij}$, which depend on the rapidities $\lambda_{j}$ and the interaction strength $c$, are determined by enforcing the cusp condition at $x_{i}=x_{j}$. This accounts for the discontinuity in the derivative due to the contact interactions [LiebLiniger]. The amplitudes are given by
@@
$\begin{split}&\lambda_{1}=J_{1}+\frac{1}{2\pi}\theta\left(\lambda_{1}-\lambda_{2}\right),\\ &\lambda_{2}=J_{2}+\frac{1}{2\pi}\theta\left(\lambda_{1}-\lambda_{2}\right)\end{split}$ (15)
where $\theta(x)=-2\arctan\left(x/c\right)$, and $J_{1}$ and $J_{2}$ are distinct relative half integers.
@@
This kind of matrix element has been obtained in Ref. [Olsen2025] for arbitrary $N$. We therefore report the explicit results for $N=2$ only. It is expressed as

### v2

@@ 4 Conclusion
Beyond its conceptual simplicity, this approach provides a powerful tool to explore Anderson localization and quantum criticality in regimes that are otherwise experimentally challenging to reach. Future directions include the development of experimental implementations using ultracold atomic gases or photonic systems, where both interactions and quasiperiodic driving can be precisely controlled. From a theoretical standpoint, our framework opens the way to investigate other symmetry classes, such as the unitary or symplectic ones, by appropriately modifying the driving protocol or introducing synthetic gauge fields. It also provides a natural platform to study properties beyond averaged quantities, including multifractality and nonergodic dynamics at criticality. Finally, while the formal mapping between the interacting Lieb-Liniger model and an effective Anderson model holds for arbitrary particle number $N$, our results explicitly demonstrate genuine Anderson localization only for $N=2$ and $N=3$ (shown in [42]). Whether increasing $N$ leads to a further growth of the effective dimensionality, or instead to a saturation of this mechanism, remains an open question that calls for future investigation.
@@ Appendix A Calculation of the matrix elements in the Bethe basis
In this section, we recall the expressions of the eigenstates of the Hamiltonian (2) from the main text for the case of $N=2$ identical bosons [33] and give the explicit expression of the matrix elements of the Floquet operator in this basis. Following the Bethe Ansatz, the wavefunction in each position sector is expressed as a superposition of plane waves with fixed energy. In the fundamental position sector $x_{1}<x_{2}$, the Bethe wavefunction takes the form
@@
where the positions $x_{j}$ are expressed in units of $\ell=L/(2\pi).$ The non-normalized amplitudes $A_{ij}$, which depend on the rapidities $\lambda_{j}$ and the interaction strength $c$, are determined by enforcing the cusp condition at $x_{i}=x_{j}$. This accounts for the discontinuity in the derivative due to the contact interactions [33]. The amplitudes are given by
@@
$\begin{split}&\lambda_{1}=I_{1}+\frac{1}{2\pi}\theta\left(\lambda_{1}-\lambda_{2}\right),\\ &\lambda_{2}=I_{2}+\frac{1}{2\pi}\theta\left(\lambda_{1}-\lambda_{2}\right)\end{split}$ (15)
where $\theta(x)=-2\arctan\left(x/c\right)$, and $I_{1}$ and $I_{2}$ are distinct relative half integers.
@@
This kind of matrix element has been obtained in Ref. [42] for arbitrary $N$. We therefore report the explicit results for $N=2$ only. It is expressed as
@@
When implemented numerically, the sum in Eq. (18) must be truncated at finite $m_{1}$ and $m_{2}$. We have verified that our results do not change if we increase these cut off parameters.